# ReviewGuard: Enhancing Deficient Peer Review Detection via LLM-Driven Data Augmentation

### v1

@@ Abstract
Peer review serves as the gatekeeper of science, yet the surge in submissions and widespread adoption of large language models (LLMs) in scholarly evaluation present unprecedented challenges. Recent work has focused on using LLMs to improve review efficiency or generate insightful review content. However, unchecked deficient reviews from both human experts and AI systems threaten to systematically undermine the peer review ecosystem and compromise academic integrity. To address this critical issue, we introduce ReviewGuard, an automated system for detecting and categorizing deficient reviews. ReviewGuard employs a comprehensive four-stage LLM-driven framework that: (1) collects ICLR and NeurIPS papers with their corresponding reviews from OpenReview; (2) annotates review types using GPT-4.1 with human validation; (3) addresses class imbalance and data scarcity through LLM-driven synthetic data augmentation, producing a final corpus of 6,634 papers, 24,657 real reviews, and 46,438 synthetic reviews; and (4) fine-tunes both encoder-based models and open source LLMs. We perform comprehensive feature analysis of the structure and quality of the review text. Compared to sufficient reviews, deficient reviews demonstrate lower rating scores, higher self-reported confidence, reduced structural complexity, and a higher proportion of negative sentiment. AI-generated text detection reveals that, since ChatGPT’s emergence, AI-generated reviews (both sufficient and deficient) have increased dramatically. In the evaluation of deficient review detection models, mixed training with synthetic and real review data provides substantial enhancements to recall and F1 scores on the binary task. For example, in Qwen 3-8B, the recall increases from 0.5499 to 0.6653, and the F1 score increases from 0.5606 to 0.7073. Similarly, for the multi-label detection task, this approach yields notable improvements in identifying deficient review types across most models. This study presents the first LLM-driven system for detecting deficient peer reviews, providing evidence to inform AI governance in peer review while offering valuable insights into human-AI collaboration to maintain academic integrity. The code, prompts, and data used in this study are publicly available at Link.
@@ I Introduction
Peer review serves as the cornerstone of scientific communication, ensuring the quality, validity, and integrity of research before publication [1, 2]. Through expert evaluation, it maintains stringent standards in methodology, precision, and scholarly contribution [3]. However, the system faces significant pressures, with submission volumes growing exponentially, overwhelming review processes, and compromising quality at scale [4]. Compounding factors include heavy reviewer workloads and compressed timelines that foster superficial evaluations, suboptimal reviewer-manuscript pairings, and reduced accountability due to anonymity [5, 6, 7, 8].
Advancements in LLMs present viable solutions, as evidenced by studies showing that LLM-generated reviews often match human assessments, pinpoint methodological problems, evaluate contributions, and provide constructive feedback [3, 2, 9, 10]. Their capacity for rapid, consistent processing of extensive text supports scalable peer review enhancements [4]. However, integrating LLMs introduces risks, such as prompt injection vulnerabilities that enable biased reviews, along with model-inherent biases, hallucinations, and generic outputs that undermine the depth and trustworthiness of the review [11, 12, 4].
Existing research on LLMs in peer review predominantly emphasizes replicating or improving review practices, with limited focus on identifying and mitigating deficient reviews. This oversight is concerning, as unchecked problematic behaviors from human or AI reviewers could exacerbate systemic flaws. Barriers to advancement include the lack of standardized datasets, metrics, and detection mechanisms, issues intensified by accelerating LLM adoption.
@@ II Related Work
Peer review is fundamental to the reliability and integrity of academic publishing, serving as both a quality control mechanism and a platform for constructive academic exchange [14, 15]. Through critical evaluation of methodological soundness, originality, and clarity, reviewers help maintain disciplinary standards and prevent the dissemination of unreliable findings [16, 17]. Reviewers’ assessments not only determine whether research enters the scientific record, but also provide authors with expert feedback that strengthens their manuscripts prior to publication [18, 19, 20]. The value of this process depends on the expertise, objectivity, and professionalism of reviewers, whose constructive and actionable feedback is essential for improving both the quality of individual manuscripts and the overall standard of scientific literature.
Despite its essential role, the peer review process continues to encounter challenges, including reviewer bias, inconsistency, limited coverage of important evaluation criteria, and an increasing burden on the research community [21]. Manual reviews are often time-consuming and costly, and numerous studies have highlighted common issues such as superficial assessments, insufficient justification, and unconstructive or careless feedback [7]. As the volume of submissions grows exponentially, these limitations have become more pronounced, raising pressing concerns about the overall quality and fairness of scientific evaluation. Recent advances in LLMs have created new opportunities to address some of these challenges. Some research has explored automated multi-agent systems that leverage LLMs to simulate and streamline the peer review process, demonstrating promising potential for improving scalability and efficiency [2, 22, 23]. Meanwhile, hierarchical and structured frameworks have been proposed to address the limitations of superficial or unsubstantiated feedback by decomposing the review process into fine-grained, aspect-specific tasks and modeling the logical relationships between papers and review comments [24, 25, 26]. Other studies have investigated augmentation strategies such as retrieval-augmented generation (RAG) and fine-tuning to further enhance review quality. RAG integrates targeted literature retrieval into the generation process to ground assessments in verifiable evidence [27, 28], while fine-tuning adapts LLMs to the stylistic and evaluative norms of peer review, improving factual accuracy and alignment with domain-specific criteria [29, 3].
@@ III Methodology
Figure 1: The ReviewGuard pipeline: A four-stage framework for automated detection of deficient peer reviews via data acquisition, human-validated and synthetic annotation, multi-dimensional feature analysis, and detection model development.
@@ III-A Peer Review Acquisition
OpenReview is a platform dedicated to enhancing transparency in scientific communication by proactively publishing both manuscripts and their corresponding peer reviews [31]. The review data from ICLR and NeurIPS are of particular value for two primary reasons. First, these conferences represent the premier venues in the field of machine learning; thus, a nuanced analysis of deficient review practices at this level has direct implications for academic integrity and scientific advancement. Second, both ICLR and NeurIPS make available a comprehensive set of peer review artifacts, including reviews for both accepted and rejected submissions, thereby ensuring greater data representativeness and mitigating selection biases often present in more restricted datasets. Furthermore, OpenReview’s official API infrastructure enables the systematic and large-scale acquisition of original manuscripts and associated peer-review texts, supporting both reproducibility and exhaustive coverage.
Although deficient review is an increasingly recognized issue, it constitutes only a minor proportion of the total review corpus. To maximize the likelihood of capturing such instances within a feasible annotation budget, our data collection is deliberately focused on manuscripts exhibiting conflicting peer review outcomes. Specifically, we operationalize “conflicting reviews” as papers where, for example, among three reviewer reports, two provide concordant, relatively high (or low) ratings while the third exhibits a substantial divergence. Formally, for each paper, we compute a consensus score by averaging all assigned reviewer scores after discarding the highest and lowest values. The absolute differences between the highest (lowest) score and the consensus, denoted as $\mathit{diff_{high}}$ and $\mathit{diff_{low}}$ respectively, quantify the extent of outlier opinions. This targeted data selection criterion enables efficient resource allocation, increasing the yield of relevant instances for subsequent annotation and in-depth empirical analysis.
@@ III-B2 Annotation Protocol
Large-scale data annotation is inherently time-consuming and labor-intensive. However, rapid advancements in LLMs have placed them as effective annotators, widely adopted in various disciplines [33]. Empirical evidence indicates that LLM-generated annotations often achieve or surpass the quality of human annotations, with inter-LLMs consistency exceeding that observed among human annotators [34, 35]. However, LLMs are not infallible oracles; their outputs require further human validation to ensure reliability. To address this, we integrate LLM-driven large-scale annotation with human verification on a smaller scale [36]. Specifically, we employ the GPT-4.1 model  [^1] GPT-4.1: https://openai.com/index/gpt-4-1/. for annotation. The input comprises the title, abstract, and full text and metadata of all associated reviews. The model is tasked with determining whether each review qualifies as sufficient; If considered deficient, it categorizes the review according to the six types outlined in Table I and provides a rationale for classification.
@@ IV-B3 Human Evaluation
To assess the accuracy and reliability of GPT-4.1’s automated annotations, we conducted a comprehensive human evaluation. The evaluation team comprised two evaluators: one undergraduate student and one graduate student, both with expertise in computer science and data science. Before formal evaluation, all team members received systematic training covering the tasks objectives, evaluation criteria, and operational procedures. To ensure objectivity and independence, communication between the evaluators was strictly prohibited during the assessment process, and all evaluations were performed independently.
In the first round of evaluation, we randomly selected 100 reviews for manual review of GPT-4.1’s annotations. The average Cohen kappa values and Fleiss kappa values for this round were: 0.4899 and 0.4740, respectively. For samples in which inconsistencies were identified in the initial assessment, the evaluators engaged in focused discussions and reached consensus through collective deliberation. Subsequently, a second round of evaluation was conducted. In this phase, another 100 reviews, distinct from the previous round, were randomly selected and the evaluation team independently assessed them using the updated consensus criteria. The average Cohen kappa values and Fleiss kappa values for this round were: 0.8774 and 0.8793, respectively. The results show that the two rounds of evaluation not only improved inter-rater consistency but also confirmed the high quality of the annotated GPT-4.1 results. Therefore, we conclude that the data automatically labeled with GPT-4.1 can serve as a reliable foundation for subsequent analyzes and experiments.
@@ IV-C1 Textual Structure Analysis
Note: Correlations shown are between the binary SR indicator (1=Sufficient Review, 0=Deficient Review) and each metric, with significance levels (*** indicates $p<0.001$). Only correlations > 0.1 are displayed. The “Mean” column shows average values for SR and DR respectively. Higher positive correlations indicate metrics that are elevated in sufficient reviews compared to deficient reviews.
@@ VI Acknowledgments
The authors acknowledge the use of LLMs in this work. Specifically, LLMs were employed for data augmentation to expand the training dataset, and for editing and grammar enhancement during the manuscript preparation process. The core research contributions, methodology, analysis, and conclusions remain the original work of the authors.

### v2

@@ Abstract
Peer review serves as the gatekeeper of science, yet the surge in submissions and widespread adoption of large language models (LLMs) in scholarly evaluation present unprecedented challenges. While recent work has focused on using LLMs to improve review efficiency, unchecked deficient reviews from both human experts and AI systems threaten to systematically undermine academic integrity. To address this issue, we introduce ReviewGuard, an automated system for detecting and categorizing deficient reviews through a four-stage LLM-driven framework: data collection from ICLR and NeurIPS on OpenReview, GPT-4.1 annotation with human validation, synthetic data augmentation yielding 6,634 papers with 24,657 real and 46,438 synthetic reviews, and fine-tuning of encoder-based models and open-source LLMs. Feature analysis reveals that deficient reviews exhibit lower rating scores, higher self-reported confidence, reduced structural complexity, and more negative sentiment than sufficient reviews. AI-generated text detection shows dramatic increases in AI-authored reviews since ChatGPT’s emergence. Mixed training with synthetic and real data substantially improves detection performance—for example, Qwen 3-8B achieves recall of 0.6653 and F1 of 0.7073, up from 0.5499 and 0.5606 respectively. This study presents the first LLM-driven system for detecting deficient peer reviews, providing evidence to inform AI governance in peer review. Code, prompts, and data are available at GitHub Repository.
@@ I Introduction
Peer review serves as the cornerstone of scientific communication, ensuring the quality, validity, and integrity of research before publication [1, 2]. Through expert evaluation, it maintains stringent standards in methodology, precision, and scholarly contribution [3]. However, the system faces significant pressures, with submission volumes growing exponentially, overwhelming review processes, and compromising quality at scale [4]. Compounding factors include heavy reviewer workloads and compressed timelines that foster superficial evaluations, suboptimal reviewer-manuscript pairings, and reduced accountability due to anonymity [5, 6, 7].
Advancements in LLMs present viable solutions, as evidenced by studies showing that LLM-generated reviews often match human assessments, pinpoint methodological problems, evaluate contributions, and provide constructive feedback [3, 2, 8]. Their capacity for rapid, consistent processing of extensive text supports scalable peer review enhancements [4]. However, integrating LLMs introduces risks, such as prompt injection vulnerabilities that enable biased reviews, along with model-inherent biases, hallucinations, and generic outputs that undermine the depth and trustworthiness of the review [9, 4, 10].
Existing research on LLMs in peer review predominantly emphasizes replicating or improving review practices, with limited focus on identifying and mitigating deficient reviews [10]. This oversight is concerning, as unchecked problematic behaviors from human or AI reviewers could exacerbate systemic flaws. Barriers to advancement include the lack of standardized datasets, metrics, and detection mechanisms, issues intensified by accelerating LLM adoption.
@@ II Related Work
Peer review is fundamental to the reliability and integrity of academic publishing, serving as both a quality control mechanism and a platform for constructive academic exchange [12, 13]. Through critical evaluation of methodological soundness, originality, and clarity, reviewers help maintain disciplinary standards and prevent the dissemination of unreliable findings [14, 15]. Reviewers’ assessments not only determine whether research enters the scientific record, but also provide authors with expert feedback that strengthens their manuscripts prior to publication. The value of this process depends on the expertise, objectivity, and professionalism of reviewers, whose constructive and actionable feedback is essential for improving both the quality of individual manuscripts and the overall standard of scientific literature.
Despite its essential role, the peer review process continues to encounter challenges, including reviewer bias, inconsistency, limited coverage of important evaluation criteria, and an increasing burden on the research community [16]. Manual reviews are often time-consuming and costly, and numerous studies have highlighted common issues such as superficial assessments, insufficient justification, and unconstructive or careless feedback [7]. As the volume of submissions grows exponentially, these limitations have become more pronounced, raising pressing concerns about the overall quality and fairness of scientific evaluation. Recent advances in LLMs have created new opportunities to address some of these challenges. Some research has explored automated multi-agent systems that leverage LLMs to simulate and streamline the peer review process, demonstrating promising potential for improving scalability and efficiency [2, 17, 18]. Meanwhile, hierarchical and structured frameworks have been proposed to address the limitations of superficial or unsubstantiated feedback by decomposing the review process into fine-grained, aspect-specific tasks and modeling the logical relationships between papers and review comments [19, 20, 21]. Other studies have investigated augmentation strategies such as retrieval-augmented generation (RAG) and fine-tuning to further enhance review quality. RAG integrates targeted literature retrieval into the generation process to ground assessments in verifiable evidence [22], while fine-tuning adapts LLMs to the stylistic and evaluative norms of peer review, improving factual accuracy and alignment with domain-specific criteria [3].
@@ III Methodology
Figure 1: The ReviewGuard pipeline for automated deficient review detection. The framework integrates four sequential stages: (1) targeted data acquisition from OpenReview focusing on conflicting reviews, (2) LLM-driven annotation with human validation and synthetic data augmentation to address class imbalance, (3) multi-dimensional feature analysis examining textual structure and quality, and (4) fine-tuning of encoder-based and open-source LLM detection models.
@@ III-A Peer Review Acquisition
OpenReview proactively publishes manuscripts and peer reviews, enhancing transparency in scientific communication [24]. ICLR and NeurIPS review data are particularly valuable because these premier machine learning conferences directly impact academic integrity, and both publish comprehensive review artifacts for accepted and rejected submissions, ensuring data representativeness and mitigating selection bias [25, 26]. OpenReview’s API enables systematic, large-scale acquisition supporting reproducibility and exhaustive coverage.
Deficient reviews constitute a minor proportion of the total corpus. To maximize detection within a feasible annotation budget, we focus on manuscripts with conflicting peer review outcomes. Conflicting reviews potentially indicate deficient reviews, as substantial divergence may reflect superficial engagement, insufficient domain expertise, or inconsistent evaluation criteria. Extreme outlier opinions may signal inadequate engagement rather than legitimate scholarly disagreement, warranting scrutiny for deficiencies in comprehension, substantiation, or constructive feedback.
We operationalize “conflicting reviews” as manuscripts where one reviewer’s assessment diverges substantially from the consensus. For each manuscript with $n$ reviews, we compute a consensus score by averaging all ratings after excluding the maximum and minimum values. The absolute differences between the extreme scores and consensus, denoted $\mathit{diff}_{\text{high}}$ and $\mathit{diff}_{\text{low}}$, quantify the extent of divergence. This selection criterion enriches the dataset with reviews more likely to exhibit deficiency patterns while maintaining representation of both sufficient and deficient reviews.
@@ III-B2 Annotation Protocol
Large-scale data annotation is inherently time-consuming and labor-intensive. However, rapid advancements in LLMs have placed them as effective annotators, widely adopted in various disciplines [28]. Empirical evidence indicates that LLM-generated annotations often achieve or surpass the quality of human annotations, with inter-LLMs consistency exceeding that observed among human annotators [29]. However, LLMs are not infallible oracles; their outputs require further human validation to ensure reliability. To address this, we integrate LLM-driven large-scale annotation with human verification on a smaller scale [30]. Specifically, we employ the GPT-4.1 model  [^1] GPT-4.1: https://openai.com/index/gpt-4-1/. for annotation. The input comprises the title, abstract, and full text and metadata of all associated reviews. The model is tasked with determining whether each review qualifies as sufficient; If considered deficient, it categorizes the review according to the six types outlined in Table I and provides a rationale for classification.
@@ IV-B3 Human Evaluation
To assess GPT-4.1’s annotation accuracy, we conducted human evaluation with two evaluators (one undergraduate and one graduate student, both with computer science and data science expertise). Evaluators received systematic training on task objectives, evaluation criteria, and procedures, then independently assessed all samples without communication to ensure objectivity.
In the first round, we randomly selected 100 reviews for manual assessment, achieving Cohen’s kappa of 0.4899 and Fleiss’ kappa of 0.4740. Analysis of inconsistent cases revealed three primary disagreement sources: (1) boundaries between “superficial” and “cursory judgment” categories, (2) thresholds distinguishing “overly harsh” from constructively critical feedback, and (3) interpretation of domain-specific terminology for “uninformed” reviews. We refined the annotation guidelines and GPT-4.1 prompt by adding concrete examples, establishing clearer decision rules, and developing a consensus framework for borderline cases.
In the second round, evaluators independently assessed another 100 reviews using updated criteria, achieving Cohen’s kappa of 0.8774 and Fleiss’ kappa of 0.8793. These results confirm substantial improvement in inter-rater agreement and validate GPT-4.1’s annotation quality, establishing the labeled data as a reliable foundation for subsequent analyses.
@@ IV-C1 Textual Structure Analysis
Note: Correlations shown are between the binary SR indicator (1=Sufficient Review, 0=Deficient Review) and each metric, with significance levels (*** indicates $p<0.001$). The “Mean” column shows average values for SR and DR respectively. Higher positive correlations indicate metrics that are elevated in sufficient reviews compared to deficient reviews.
@@ VI Acknowledgments
This research is partially supported by NSF grant #2225229. The authors acknowledge the use of LLMs in this work. Specifically, LLMs were employed for data augmentation to expand the training dataset, and for editing and grammar enhancement during the manuscript preparation process. The core research contributions, methodology, analysis, and conclusions remain the original work of the authors.